At the city museum, child admission is $5.30 and adult admission is $9.10 On Wednesday, 163 tickets were sold for a total sales of $1198.30 How many child tickets were sold that day?
Answer by jorel1380(3719)

p+q=163
5.3p+9.1q=1198.3
53p+91q=11983
53p+53q=8639
38q=3344
q=88
p=75
75 childrens tickets were sold that day..